OpenStudy (anonymous):

6=2(x+8)-5x WILL GIVE MEDAL FOR CORRECT ANSWER!

OpenStudy (anonymous):

10/3

OpenStudy (anonymous):

6=2x+16-5x combine like terms 6=-3x +16 -10=-3x 10/3=x